# ConsistentEE: A Consistent and Hardness-Guided Early Exiting Method for Accelerating Language Models Inference

### 19 Dec 2023 v1

@@ Abstract
Early Exiting is one of the most popular methods to achieve efficient inference. Current early exiting methods adopt the (weighted) sum of the cross entropy loss of all internal classifiers during training, imposing all these classifiers to predict all instances correctly. However, during inference, as long as one internal classifier predicts an instance correctly, it can accelerate without losing accuracy. Thus, there is a notable gap between training and inference. We propose ConsistentEE, an early exiting method that is consistent in training and inference. ConsistentEE formulates the early exiting process as a reinforcement learning problem. A policy network is added to decide whether an instance should exit or continue. The training objective of ConsistentEE only require each instance to be predicted correctly by one internal classifier. Additionally, we introduce the concept Memorize Layer to measure the hardness of an instance. We incorporate memorized layer into reward function design, which allows “easy” instances to focus more on acceleration while “hard” instances to focus more on accuracy. Experimental results show that our method outperforms other baselines on various natural language understanding and generation tasks.
@@ Introduction
As one of the most popular methods, early exiting adds an internal classifier to each intermediate layer, allowing instances to stop model inference in an early layer instead of going though the entire model, thus accelerating the inference time. Early Exiting can be useful especially in the era of LLMs. Although the scale of LLMs continues to increase to deal with challenging tasks like reasoning, planning, etc., in the real-world scenario, they may not always encounter such challenging tasks (Schuster et al. 2022). A smaller sized LLM has been good at handle “easy” tasks such as paraphrasing, translation, etc. When LLMs encounter “easy” tasks, early exiting can avoid going though the entire model, thus saving inference time.
@@
Current early exiting methods typically adopt the (weighted) sum of the cross entropy (CE) loss of all internal classifiers as the training loss, which imposes that all internal classifiers should predict all instances correctly. However, our experimental analysis (in Figure 2 left) shows that the classification accuracy of each internal classifier is not so satisfactory as expected. This observation raises a question: is it necessary to require all internal classifiers to predict each instance correctly? Actually, in the inference phase, as long as one internal classifier predicts an instance correctly, the process can be accelerated without sacrificing accuracy. Given that there is no such strict requirement during inference, it is reasonable to abandon such a demanding objective during the training phase.
@@
We propose ConsistentEE, an early exiting method that is consistent in training and inference. Specifically, ConsistentEE formulates the training process as a reinforcement learning (RL) problem. As shown in Figure 1, a policy network is added at each intermediate layer to determine whether to exit or not. If a policy network at an intermediate layer decides to exit, the corresponding internal classifier is expected to predict this instance correctly, otherwise no loss is incurred by this classifier. During training, an instance can exit at only one layer. Hence, only one internal classifier is required to predict it correctly. Our experimental analysis, as depicted in Figure 2 right, reveals that unlike existing methods, the ConsistentEE objective enables each layer’s classification accuracy to consistently remain at a high level.
@@
For the reward function design, we consider both accuracy and acceleration. To ensure accuracy, we incorporate the loss of the internal classifier into the reward function, assigning higher reward values to instances with lower losses. Additionally, to consider acceleration, we involve the layer depth at which an instance exits into the reward function. Although one can place a trade-off coefficient to balance accuracy and acceleration in the reward function, we argue that instances of different hardness levels should put different weights on accuracy and acceleration. We observe from experimental results that “easy” instances generally can be classified correctly at shallow layers. Those instances should exit as early as possible after it can be classified correctly. “Hard” instances typically can be classified correctly at deep layers. Those instances should focus on accuracy instead of acceleration at early layers. Hence, we also incorporate the hardness level of an instance into the reward function design.
@@
However, the identification of “easy” and “hard” instances is itself a difficult problem and is extensively studied in the literature (Kumar, Packer, and Koller 2010; Arpit et al. 2017; Toneva et al. 2019). Inspired by the concept of unforgettable example (Toneva et al. 2019), we propose a new concept, Memorized Layer, to measure the hardness. Memorized layer is the layer where the instance is correctly classified and continuingly correctly classified until the final layer. Experimental analysis reveals a high and medium correlation between the memorized layer and losses, and forgetting events respectively. Our reward function incorporates “memorized layer” to encourage “easy instances” to pay more attention to acceleration while “hard instances” to focus on accuracy rather than acceleration.
@@
Confidence-based early exiting methods utilize confidence, entropy, or (calibrated) max class probability to exit. DeeBERT (Xin et al. 2020) trained the internal classifiers by minimizing the sum of their cross-entropy losses. During inference, the instance exits if the entropy is less than a predefined threshold. FastBERT (Liu et al. 2020) distilled the final classifier layer to all internal classifiers and used entropy as the exiting criterion. The training objective of FastBERT is the sum of distillation losses. Other early exiting methods such as RomeBERT (Geng et al. 2021) and Past-Future (Liao et al. 2021) also used entropy as the exiting criterion. Right-Tool (Schwartz et al. 2020), SkipBERT (Wang et al. 2022), and CascadeBERT (Li et al. 2021) use (calibrated) max class probability as the exiting criterion.
Ensemble-based early exiting methods utilize predictions from multiple internal classifiers to make better decisions. PABEE (Zhou et al. 2020) adopted a weighted sum of cross-entropy losses as the training objective. During inference, the instance exits when $k$ consecutive internal classifiers make the same prediction. PCEE-BERT (Zhang et al. 2022) combined both ensemble-based and confidence-based exiting criteria. The instance exits if the confidence scores are greater than a predefined threshold for several consecutive exits.
@@
The above methods are early exiting encoding methods to accelerate BERT inference. CALM (Schuster et al. 2022) focused on token-wise early exit decoding strategy to achieve efficient inference on large language models. CALM proposed three confidence measures as exiting criteria including Softmax response, hidden-state saturation, and early exit classifier. A calibration procedure is proposed to find a shared exit threshold.
@@
During training, this loss (Eq. 2) imposes that all internal classifiers should predict all instances correctly. To investigate whether such an objective is suitable, we analyze the loss values and accuracy of each internal classifier on the RTE dataset. Additional results of some other datasets are included in the Appendix. As shown in Figure 2 (left), with the traditional training loss, the internal classifiers do not perform as expected. The accuracy of shallow layers is unsatisfactory. Especially, the first three layers fall below the random guess baseline, i.e., 50%. We think that the traditional objective is too strict that the requirement may have exceeded the limits of the capabilities of internal classifiers, thus harming their performance.
The experimental analysis provokes a question: whether it is necessary to require all internal classifiers to predict each instance correctly? Actually, in the inference phase, as long as one internal classifier predicts an instance correctly, it can accelerate without sacrificing accuracy. We therefore propose ConsistentEE, an early exiting method that is consistent in training and inference. With ConsistentEE, for a training instance, only one internal classifier is required to predict it correctly and this instance will exit at the corresponding layer. ConsistentEE allows each classifier to focus on correctly classifying partial instances, reducing the burden of each layer. As shown in Figure 2, the training losses under ConsistentEE are consistently lower than those under the traditional training objective at every layer. Additionally, the accuracy of each layer in the training set remains at a significantly high level.
@@
The primary challenge in ConsistentEE is to determine the appropriate layer at which an instance should exit. As the ground truth exit layer is unavailable, ConsistentEE employs the reinforcement learning (RL) method to automatically learn the optimal layer for an instance to exit during training. As shown in Figure 1, a policy network is added to an intermediate layer besides an internal classifier. This policy network shares the same input with the internal classifier. It produces the probability distribution of two actions. One is exiting at the current layer, and the other one is continuing to the next layer. If the policy network takes an action to exit, then the internal classifier at the same layer is required to predict the instance correctly. Otherwise, if the policy network takes an action to continue, then there is no loss imposed on the corresponding internal classifier.
@@
Early exiting can be regarded as a sparse reward process because rewards are only received when an instance decides to exit at a particular layer, with no feedback provided for intermediate layers. Considering this characteristic, we use the Policy Gradient technique to automatically discover the optimal exit layer for each instance. We introduce relevant notations in the following.
State $s_{t}$ is the representation of the input at $t$-th layer. For classification task with PLMs as backbones, $s_{t}$ is the representation of [CLS] token at $t$-th layer. For generation task with decoder-only LLMs as backbones, $s_{t}$ is the hidden state of the last token of the input at $t$-th layer. The hidden state is the output of the last layer (i.e., Add & Norm layer) in the basic Transformer block.
@@
In the vanilla version of the reward function, each instance has the same trade-off coefficient to balance the accuracy and acceleration. However, “easy” instances generally can be correctly classified at shallow layers. Such instances should exit as early as possible after it can be correctly classified. “Hard” instances typically can be classified correctly at deep layers. Such instances should focus on accuracy instead of acceleration at early layers. Hence, instances of different hardness levels should put different weights on accuracy and acceleration. Based on this idea, we improve the reward function by taking instance hardness into account:
@@
where $L$ is the total number of layers; ${\mathcal{M}}$ denotes Memorized Layer, a new concept we introduce to measure the hardness of an instance. A smaller (larger) value of ${\mathcal{M}}$ indicates an easier (harder) instance. More details about this concept will be introduced in the next section. With Eq. 4, for a hard (easy) instance, the trade-off coefficient $\alpha\cdot(1-\frac{{\mathcal{M}}}{L})$ is expected to be smaller (larger), causing it to receive more (less) weight on accuracy compared to the vanilla version.
@@
Task Objective Function We optimize the internal classifiers and the backbone to accomplish the task (e.g., classification, generation etc.) The internal classifiers and the backbone are parameterized by $\omega$. The task objective function is defined as:
@@
where $\mathds{1}(\cdot)$ is an indicator function that returns 1 if the condition is satisfied and 0 otherwise.
@@
To evaluate acceleration capacities on classification task with PLMs as backbones, we conduct experiments on six classification datasets of the GLUE benchmark (Wang et al. 2019) and two multi-classes classification datasets including M-CID (Arora et al. 2020) and StackOverflow (Xu et al. 2015). M-CID and StackOverflow are intent classification datasets which are collected from Covid-19 services and online questions. Since most GLUE datasets are binary classification datasets, it is possible that instances may still be classified correctly despite experiencing premature or hasty exits. Hence we evaluate early exiting methods on some multi-classes classification datasets. Statistics of these datasets are listed in Table 2. We compared our method with various baselines including DeeBERT(Xin et al. 2020), PABEE (Zhou et al. 2020), BERxiT (Xin et al. 2021), Right-Tool (Schwartz et al. 2020), PCEE-BERT (Zhang et al. 2022), HashEE (Sun et al. 2022), and TR-BERT (Ye et al. 2021). For layer-wise exiting methods, we use relatively saved layers to evaluate model acceleration. According to (Xin et al. 2020), saved layers are linearly proportional to actual runtime. While for token-wise exiting methods, we use relative saved runtime. Most results are reported on the development set, since the large number of evaluations are restricted by the GLUE evaluation server. BERT-Base (Devlin et al. 2019), two competitive baselines, and ConsistentEE are evaluated on the test set.
To evaluate acceleration capacities on classification task with LLMs as backbones, we perform the supervised fine-tuning step with LoRA (Hu et al. 2022) using Alpaca (Taori et al. 2023) as training data and test on the Dolly dataset (Conover et al. 2023). The backbone LLMs are LLaMA-7b and LLaMA-13b. We set the beam to 1, top-sampling to 0.75, top-k selection to 40 and temperature to 1. The baseline is CALM (Schuster et al. 2022). We use Rouge-L (Lin 2004) and BERT-F (Zhang et al. 2019) scores to measure the model quality and use save layers to measure the model acceleration. Following (Elbayad et al. 2020; Schuster et al. 2022), we use a state copying strategy to handle the computation of hidden states. Specifically, the computation of the input hidden state of token $t$ at layer $i$ relies on the output hidden states of the previous layer $i-1$ for all preceding tokens $\{1,\cdots,t-1\}$. In cases where some preceding tokens exit before the $i-1$ layer, we copy the output hidden state at the layer at which the token exits, and use it as the output hidden state in $i-1$ layer.
@@
As depicted in Figure 4, the performance of the vanilla reward and the hardness-guided reward are comparable when the speedup ratio is small. However, as the speedup ratio increases, the performance of the vanilla reward deteriorates significantly, while the hardness-guided reward maintains a satisfactory level of accuracy. The possible reason is that the vanilla reward treats each instance equally, causing hard instances struggle to maintain accuracy as they prioritize acceleration similar to easy instances. In contrast, the hardness-guided reward encourages instances to assign different weights to accuracy and acceleration based on their individual hardness. Simple samples can be correctly classified effortlessly and achieve as much acceleration as possible. On the other hand, hard instances place a greater emphasis on accuracy while paying less attention to acceleration, which is more likely to achieve a satisfactory level of accuracy. This differentiated treatment allows the hardness-guided reward to effectively balance the trade-off between accuracy and acceleration across various instances.
@@
As shown in Figure 6, ConsistentEE and CALM demonstrate similar performance when the speedup ratio is below 2x. However, as the speedup ratio increases, ConsistentEE outperforms CALM. Examples of generation from ConsistentEE under different speedup ratios result are shown Table 4. The responses generated under 2.0x appear reasonable and closely resemble the original response. However, the responses generated under 5.0x vary from the original response but still make sense.

### 07 Apr 2024 v2

@@ Abstract
Early Exiting is one of the most popular methods to achieve efficient inference. Current early exiting methods adopt the (weighted) sum of the cross entropy loss of all internal classifiers as the objective function during training, imposing all these classifiers to predict all instances correctly. However, during inference, as long as one internal classifier predicts an instance correctly, it can accelerate without losing accuracy. Thus, there is a notable gap between training and inference. We propose ConsistentEE, an early exiting method that is consistent in training and inference. ConsistentEE formulates the early exiting process as a reinforcement learning problem. A policy network is added to decide whether an instance should exit or continue. The training objective of ConsistentEE only requires each instance to be predicted correctly by one internal classifier. Additionally, we introduce the concept Memorized Layer to measure the hardness of an instance. We incorporate the memorized layer into reward function design, which allows “easy” instances to focus more on acceleration while “hard” instances to focus more on accuracy. Experimental results show that our method outperforms other baselines on various natural language understanding and generation tasks using PLMs and LLMs as backbones respectively.
@@ Introduction
As one of the most popular methods, early exiting adds an internal classifier to each intermediate layer, allowing instances to stop model inference in an early layer instead of going through the entire model, thus accelerating the inference time. Early Exiting can be useful, especially in the era of LLMs. Although the scale of LLMs continues to increase to deal with challenging tasks like reasoning, planning, etc., in the real-world scenario, they may not always encounter such challenging tasks (Schuster et al. 2022). A small sized LLM has been good at handling “easy” tasks such as paraphrasing, translation, etc. When LLMs encounter “easy” tasks, early exiting can avoid going through the entire model, thus saving inference time.
@@
Current early exiting methods typically adopt the (weighted) sum of the cross entropy (CE) loss of all internal classifiers as the training loss, which imposes that all internal classifiers should predict all instances correctly. However, our experimental analysis (in Figure 7 left) shows that the classification accuracy of each internal classifier is not as satisfactory as expected. This observation raises a question: is it necessary to require all internal classifiers to predict each instance correctly? Actually, in the inference phase, as long as one internal classifier predicts an instance correctly, the inference can be accelerated without sacrificing accuracy. Given that there is no such strict requirement during inference, it is reasonable to abandon such a demanding objective during the training phase.
@@
We propose ConsistentEE, an early exiting method that is consistent in training and inference. Specifically, ConsistentEE formulates the training process as a reinforcement learning (RL) problem. As shown in Figure 1, a policy network is introduced at each intermediate layer to determine whether to exit or not. If a policy network at an intermediate layer decides to exit, the corresponding internal classifier is expected to predict this instance correctly, otherwise no loss is incurred by this classifier. During training, an instance can exit at only one layer. Hence, only one internal classifier is required to predict it correctly. As depicted in Figure 7 right, unlike existing methods, the ConsistentEE objective enables each layer’s classification accuracy to consistently remain at a high level.
@@
For the reward function design, we consider both accuracy and acceleration. To ensure accuracy, we incorporate the loss of the internal classifier into the reward function, assigning higher reward values to instances with lower losses. Additionally, to consider acceleration, we involve the layer depth at which an instance exits into the reward function. Although one can place a trade-off coefficient to balance accuracy and acceleration in the reward function, we argue that instances of different hardness levels should put different weights on accuracy and acceleration. We observe that “easy” instances generally can be classified correctly at shallow layers. Those instances should exit as early as possible after they can be classified correctly. “Hard” instances typically can be classified correctly at deep layers. Those instances should focus on accuracy instead of acceleration at early layers. Hence, we also incorporate the hardness level of an instance into the reward function design.
@@
However, the identification of “easy” and “hard” instances is itself a difficult problem and is extensively studied in the literature (Kumar, Packer, and Koller 2010; Arpit et al. 2017; Toneva et al. 2019). Inspired by the concept of unforgettable example (Toneva et al. 2019), we propose a new concept, Memorized Layer, to measure the hardness. The memorized layer is the layer where the instance is correctly classified and continuously correctly classified until the final layer. Experimental analysis reveals a high and medium correlation between the memorized layer and losses, and forgetting events respectively. Our reward function incorporates the “memorized layer” to encourage “easy instances” to pay more attention to acceleration while “hard instances” to focus on accuracy rather than acceleration.
@@
Confidence-based early exiting methods utilize confidence, entropy, or (calibrated) max class probability to exit. In DeeBERT (Xin et al. 2020), FastBERT (Liu et al. 2020), RomeBERT (Geng et al. 2021) and Past-Future (Liao et al. 2021), the instance exits if the entropy is less than a predefined threshold. Right-Tool (Schwartz et al. 2020), SkipBERT (Wang et al. 2022), and CascadeBERT (Li et al. 2021) use (calibrated) max class probability as the exiting criterion.
Ensemble-based early exiting methods utilize predictions from multiple internal classifiers to make better decisions. In PABEE (Zhou et al. 2020), the instance exits when $k$ consecutive internal classifiers make the same prediction. PCEE-BERT (Zhang et al. 2022) combined both ensemble-based and confidence-based exiting criteria. The instance exits if the confidence scores are greater than a predefined threshold for several consecutive exits.
@@
The above methods are early exiting encoding methods to accelerate BERT inference. CALM (Schuster et al. 2022) and Free (Bae et al. 2023) focused on early-exiting for autoregressive models that allocates adaptive computation paths for each token generation. CALM proposed three confidence measures as exiting criteria including Softmax response, hidden-state saturation, and early exit classifier. A calibration procedure is proposed to find a shared exit threshold. Free (Bae et al. 2023) is concurrent with our work. FREE only allowed two exit points and replaced the state copying mechanism with synchronized parallel decoding to prevent performance degradation.
@@
During training, this loss (Eq. 2) imposes that all internal classifiers should predict all instances correctly. To investigate whether such an objective is suitable, we analyze the loss values and accuracy of each internal classifier on the RTE dataset. As shown in Figure 7 (left), with the traditional training loss, the internal classifiers do not perform as expected. The accuracy of shallow layers is unsatisfactory. Especially, the first three layers fall below the random guess baseline, i.e., 50%. We think that the traditional objective is too strict that the requirement may have exceeded the limits of the capabilities of internal classifiers, thus harming their performance.
The experimental analysis provokes a question: whether it is necessary to require all internal classifiers to predict each instance correctly? Actually, in the inference phase, as long as one internal classifier predicts an instance correctly, it can accelerate without sacrificing accuracy. We therefore propose ConsistentEE, an early exiting method that is consistent in training and inference. With ConsistentEE, for a training instance, only one internal classifier is required to predict it correctly and this instance will exit at the corresponding layer. ConsistentEE allows each classifier to focus on correctly classifying partial instances, reducing the burden of each layer. As shown in Figure 7, the training losses under ConsistentEE are consistently lower than those under the traditional training objective at every layer. Additionally, the accuracy of each layer in the training set remains at a significantly high level. The accuracy of layer $1$ to $4$ is not available becauase if the policy network chooses not to exit at these layers, then there is no loss and accuracy in these layers. RTE is a challenging dataset, and no samples exit at these layers.
@@
The primary challenge in ConsistentEE is to determine the most appropriate layer for an instance to exit. As the ground truth exit layer is unavailable, ConsistentEE employs the reinforcement learning (RL) method to automatically learn the optimal layer for an instance to exit during training. As shown in Figure 1, a policy network is introduced to an intermediate layer besides an internal classifier. This policy network shares the same input with the internal classifier. It produces the probability distribution of two actions. One is exiting at the current layer, and the other one is continuing to the next layer. If the policy network takes an action to exit, then the internal classifier at the same layer is required to predict the instance correctly. Otherwise, if the policy network takes an action to continue, then there is no loss imposed on the corresponding internal classifier.
@@
Early exiting can be regarded as a sparse reward process because rewards are only received when an instance decides to exit at a particular layer, with no feedback provided for intermediate actions. Considering this characteristic, we use the Policy Gradient technique to automatically discover the optimal exit layer for each instance. We introduce relevant notations in the following.
State $s_{t}$ is the representation of the input at $t$-th layer. For the classification task with PLMs as backbones, $s_{t}$ is the representation of [CLS] token at $t$-th layer. For the generation task with decoder-only LLMs as backbones, $s_{t}$ is the hidden state of the last token of the input at $t$-th layer. The hidden state is the output of the last layer (i.e., Add & Norm layer) in the basic Transformer block.
@@
In the vanilla version of the reward function, each instance has the same trade-off coefficient to balance the accuracy and acceleration. However, “easy” instances generally can be correctly classified at shallow layers. Such instances should exit as early as possible after they can be correctly classified. “Hard” instances typically can be classified correctly at deep layers. Such instances should focus on accuracy instead of acceleration at early layers. Hence, instances of different hardness levels should put different weights on accuracy and acceleration. Based on this idea, we improve the reward function by taking instance hardness into account:
@@
where $L$ is the total number of layers; ${\mathcal{M}}$ denotes Memorized Layer, a new concept we introduced to measure the hardness of an instance. A smaller (larger) value of ${\mathcal{M}}$ indicates an easier (harder) instance. More details about this concept will be introduced in the next section. With Eq. 4, for a hard (easy) instance, the trade-off coefficient $\alpha\cdot(1-\frac{{\mathcal{M}}}{L})$ is expected to be smaller (larger), causing it to receive more (less) weight on accuracy compared to the vanilla version.
@@
Task Objective Function We optimize the internal classifiers and the backbone to accomplish the task (e.g., classification, generation, etc.) The internal classifiers and the backbone are parameterized by $\omega$. The task objective function is defined as:
@@
where $\mathds{1}(\cdot)$ is an indicator function that returns $1$ if the condition is satisfied and 00 otherwise.
@@
To evaluate acceleration capacities on the classification task with PLMs as backbones, we conduct experiments on six classification datasets of the GLUE benchmark (Wang et al. 2019) and two multi-classes classification datasets including M-CID (Arora et al. 2020) and StackOverflow (Xu et al. 2015). M-CID and StackOverflow are intent classification datasets which are collected from Covid-19 services and online questions. Since most GLUE datasets are binary classification datasets, it is possible that instances may still be classified correctly despite experiencing premature or hasty exits. Hence we evaluate early exiting methods on some multi-classes classification datasets. Statistics of these datasets are listed in Table 2. We compared our method with various baselines including DeeBERT(Xin et al. 2020), PABEE (Zhou et al. 2020), BERxiT (Xin et al. 2021), Right-Tool (Schwartz et al. 2020), PCEE-BERT (Zhang et al. 2022), HashEE (Sun et al. 2022), and TR-BERT (Ye et al. 2021). For layer-wise exiting methods, we use saved layers to evaluate model acceleration. According to (Xin et al. 2020), saved layers are linearly proportional to actual runtime. While for token-wise exiting methods, we use saved runtime. Most results are reported on the development set, since the large number of evaluations are restricted by the GLUE evaluation server. Only BERT-Base (Devlin et al. 2019), three competitive baselines, and ConsistentEE are evaluated on the test set.
To evaluate acceleration capacities on the generation task with LLMs as backbones, we perform the supervised fine-tuning step using Alpaca (Taori et al. 2023) as training data and test on the Dolly dataset (Conover et al. 2023). We also evaluate our method on text summarization task on CNN/DM dataset (Nallapati et al. 2016). The backbone LLMs are LLaMA-7b and LLaMA-13b (Touvron et al. 2023). Due to our limited computational resources, we train the LLM backbone with LoRA (Hu et al. 2022) in the initialization step. Subsequently, we freeze the LLM backbone and only train the internal classifiers and policy networks. To save pre-processing time, we adopt the vanilla reward function rather than the hardness-guided reward function because the time complexity of computing the memorized layer is linear to the number of tokens in generated responses. We set the beam to 1, top-sampling to 0.75, top-k selection to 40 and temperature to 1. The baseline is CALM (Schuster et al. 2022). We use Rouge-L (Lin 2004) and BERT-F (Zhang et al. 2019) scores to measure the model quality and use save layers to measure the model acceleration. When we compute Rouge-L and BERT-F, we use the response without acceleration as the reference on Alpaca/Dolly dataset and use the gold summary as the reference on CNN/DM. Following (Elbayad et al. 2020; Schuster et al. 2022), we use a state copying strategy to handle the computation of hidden states. Specifically, the computation of the input hidden state of token $t$ at layer $i$ relies on the output hidden states of the previous layer $i-1$ for all preceding tokens $\{1,\cdots,t-1\}$. In cases where some preceding tokens exit before the $i-1$ layer, we copy the output hidden state at the layer at which the token exits, and use it as the output hidden state in $i-1$ layer.
@@
As depicted in Figure 4, the performance of the vanilla reward and the hardness-guided reward are comparable when the speedup ratio is small. However, as the speedup ratio increases, the performance of the vanilla reward deteriorates significantly, while the hardness-guided reward maintains a satisfactory level of accuracy. The possible reason is that the vanilla reward treats each instance equally, causing hard instances struggle to maintain accuracy as they prioritize acceleration as easy instances do. In contrast, the hardness-guided reward encourages instances to assign different weights to accuracy and acceleration based on their individual hardness. Simple samples can be correctly classified effortlessly and achieve as much acceleration as possible. On the other hand, hard instances place a greater emphasis on accuracy while paying less attention to acceleration, which is more likely to achieve a satisfactory level of accuracy. This differentiated treatment allows the hardness-guided reward to effectively balance the trade-off between accuracy and acceleration across various instances.
@@
As shown in Figure 6, on the Alpaca/Dolly dataset, ConsistentEE and CALM demonstrate similar performance when the speedup ratio is below 2x. However, as the speedup ratio increases, ConsistentEE outperforms CALM. On the CNN/DM dataset, ConsistentEE outperforms CALM consistently under different speedup ratios. Examples of generation from ConsistentEE under different speedup ratios result are shown in Table 5 in Appendix. The responses generated under 2.0x appear reasonable and closely resemble the original response. However, the responses generated under 5.0x vary from the original response but still make sense.
Conclusion
We propose a reinforcement learning based approach to early exiting, so that at the training phase, only one internal classifier is required to predict the instance correctly. This makes the training phase consistent with the inference phase and can allow each layer to obtain better classification accuracy. For the reward function of the reinforcement learning framework, we propose the concept memorized layer to measure the hardness of each instance, and use it to dynamically balance accuracy and acceleration instead of using a fixed coefficient. Experimental results conducted on various datasets show that our approach is able to outperform the competitive baselines, demonstrating its effectiveness and efficiency.
Acknowledgement
Appendix
A: Loss Values
In this part, We show more results about loss values on training and development sets on RTE, and MRPC datasets. In Figure 7 (a) - (d) left, we found that the loss values under ConsistentEE are lower than that under the traditional loss, and the accuracy of each layer under the ConsistentEE objective is consistently higher than that under the traditional objective. This shows that a harsh training objective does not lead to an ambitious goal where all internal classifiers should predict all instances correctly. In Figure 7 (a) - (d) right, we found that the pattern of the loss values and accuracy on the development set is similar to that on the training set. The accuracy of each layer under both objectives decreases on the development set. The accuracy under the traditional objective is stable after a specific layer. The accuracy under the ConsistentEE decreases after a specific layer because a “hard” instance is more likely to exit at a deeper layer. Deeper classifiers encounter many such “hard” instances.
(a) Traditional Objectives on RTE dataset.
(b) ConsistentEE Objectives on RTE dataset.
(c) Traditional Objectives on MRPC dataset.
(d) ConsistentEE Objectives on MRPC dataset.
Figure 7: Loss values at different layers on training and development dataset using the weighted sum objective and ConsistentEE objective respectively. The dashed dot line is the classification boundary. A loss above the boundary means misclassification. The dashed line is the classification accuracy of each layer. The darker the color, the more samples share the same loss value.
B: Results on different PLM backbones:
We provide more results on different PLM backbones including ALBERT-Base, BERT-Large, ALBERT-Large in Table 4. Our method consistently outperforms baselines.
Method RTE MRPC SST-2
Acc
Layer
F1
Layer
Acc
Layer
ALBERT-Base
75.1
12
86.7
12
94.1
12
PABEE
74.4
-10%
86.3
-28%
93.7
-37%
PCEE-BERT
74.7
-18%
86.5
-23%
93.9
-38%
ConsistentEE 75.0 -25% 86.7 -34% 94.0 -43%
BERT-Large
70.1
24
89.3
24
94.8
24
PABEE
69.0
-19%
86.6
-13%
93.8
-22%
PCEE-BERT
69.3
-17%
88.9
-12%
93.8
-25%
ConsistentEE 69.5 -26% 89.2 -22% 94.1 -30%
ALBERT-Large
83.1
24
92.0
24
94.9
24
PABEE
81.2
-15%
91.5
-26%
94.8
-22%
PCEE-BERT
82.2
-18%
91.6
-36%
94.5
-27%
ConsistentEE 82.5 -24% 91.8 -39% 94.8 -35%
Table 4: F1 / Accuracy on RTE, MRPC, SST-2 test sets.
C: Case study on easy and hard examples
Examples on Classification task (SST-2) on the backbone of BERT-base (12 layers):
Easy (exiting at 2nd layer): (1) “it ’s a charming and often affecting journey;” (2)“i had to look away - this was god awful;”
Hard (exiting at 12th layer): (1)“pumpkin takes an admirable look at the hypocrisy of political correctness , but it does so with such an uneven tone that you never know when humor ends and tragedy begins; ” (2)“this is the sort of burly action flick where one coincidence pummels another , narrative necessity is a drunken roundhouse , and whatever passes for logic is a factor of the last plot device left standing . ”
Examples on Generation task (Alpaca/Dolly) on the backbone of LLaMA-7b (32 layers):
Easy (averagely exiting at 7.73th layer): (1) Instruction: What is Python? Response: A computer programming language that can impart APIs, third-party apps, and web content. (2) Instruction: Why mobile is bad for human? Response: The use of mobile phones has been found to have a negative impact on human well-being. The tracking of human activities through mobile phone usage has been shown to reduce cognitive processes and increase the level of social limitations.
Hard (averagely exiting at 31.17th layer): (1) Instruction: What is a dispersive prism? Response: A dispersive prism is an optical prism that is used to disperse light, that is, to separate light into its spectral components (the colors of the rainbow). Different wavelengths (colors) of light will be deflected by the prism at different angles. (2) Instruction: write a scene between two actors discussing movie Inception. Response: Actor 1: ”Have you seen Inception?” Actor 2: ”Yes, I have. It was an amazing movie.” Actor 1: ”I know, right? I was so impressed with the visuals and the storyline.” Actor 2: ”Yeah, I agree. It was a great movie.” Actor 1: ”I’m so glad I saw it. It was so thought-provoking.” Actor 2: ”I know, it really made me think about the concept of dreams and reality.”
D: Examples of generation from ConsistentEE
Examples of generation from ConsistentEE under different speedup ratios result on Alpaca/Dolly are shown in Table 5 and Table 6. The responses generated under 2.0x appear reasonable and closely resemble the original response. However, the responses generated under 5.0x vary from the original response but still make sense.